OpenStudy (anonymous):

is x^2-6x-7=0 equal 2 x=-1 or x=7

OpenStudy (anonymous):

x^2 - 6x- 7 = 0 x^2 +x - 7x - 7 = 0 x(x+1) - 7(x+1) = 0 (x-7)(x+1) = 0 hence x = -1 or x=7 u r right

OpenStudy (anonymous):

i hope u have got a good answer

OpenStudy (anonymous):

Yeah i did, thanks (just wasnt sure) #ya know ?

OpenStudy (anonymous):

give me medal then

OpenStudy (anonymous):

cooochie ?

OpenStudy (anonymous):

what does coochie means